# SSHPool: The Separated Subgraph-based Hierarchical Pooling

### v1

@@ IV-A Comparisons with Classical GNN and Pooling Methods
Experimental Settings: We evaluate the classification performance of our GNN-based framework associated with the proposed SSHPool operation, and compare it against several alternative methods, including one classical graph kernel associated with C-Support Vector Machine (C-SVM) [34], three GNN models, and four GNN-models associated with different graph pooling operations. All these methods for the evaluation are introduced as follows.
@@
  • SUGAR [19] is an embedding-based GNN model with hierarchical subgraph-level selection.
  • DiffPool [17] is a GNN model with differentiable graph pooling module that generated representations of coarsened graphs. In the hierarchical architecture, there is an assignment matrix to compress the graph.
@@
For the alternative graph kernels, we search the optimal hyperparameters for the graph kernel on each dataset. For the GNN models or the GNN models associated with the proposed and alternative graph pooling operations, the Adam optimizer is set with the same learning rate of $45e-$ 4 in our experiments, and each method is trained with 20 epochs. Moreover, during the training process of these GNN models, we set the hidden dimension as 20, the dropout as 0.5, and the batch size as 16. In order to compare all these methods fairly, for either the graph kernels or the GNN models we adopt the 10-fold cross validation to compute the classification accuracy, and repeat the experiment for 10 times on each dataset. We show the average classification accuracy ($\pm$ standard error) and the averaged rank for each method in Table II. Finally, for the proposed SSHPool, we set the largest layer $L$ as $3$, then node numbers of the three layers follow $\{128,64,32\}$.
@@
Datasets D&D PROTEINS ENZYMES IMDB-B IMDB-M Avg. Rank
WLSK 74.38$\pm$0.69 71.70$\pm$0.67 25.23$\pm$0.28 64.48$\pm$0.90 43.38 $\pm$0.75 9.8
DGCNN 75.34$\pm$0.78 73.21$\pm$0.34 44.30$\pm$0.93 67.45$\pm$0.83 46.33 $\pm$0.73 8.6
SUGAR 78.62$\pm$0.26 76.81$\pm$0.31 41.56$\pm$0.74 71.43$\pm$0.93 48.12$\pm$1.76 5.6
DiffPool 77.24$\pm$0.46 74.86$\pm$0.35 46.67$\pm$1.75 70.12$\pm$0.63 47.20$\pm$1.81 7
ASAPool 75.75$\pm$0.45 75.53$\pm$0.94 49.80$\pm$1.06 71.25$\pm$0.58 48.67$\pm$1.34 5.2
ABDPool 74.13$\pm$0.52 73.24$\pm$0.91 45.24$\pm$0.96 70.58$\pm$0.71 50.63$\pm$1.47 6.6
SEPool 77.43$\pm$0.23 75.32$\pm$0.34 53.05$\pm$1.76 70.34$\pm$0.53  50.31$\pm$1.72 4.4
AdamGNN 80.12$\pm$0.34 77.04$\pm$0.78 50.72$\pm$0.35 72.85$\pm$0.67 49.35$\pm$0.96 2.8
SSHPool(non) 79.31$\pm$0.36 77.18$\pm$0.43 47.39$\pm$0.79 71.80$\pm$0.39 48.13$\pm$1.03 3.8
SSHPool 81.80$\pm$0.48 79.38$\pm$0.28 51.31$\pm$1.34 73.02$\pm$0.47 51.14$\pm$1.28 1.2
@@
Experimental Results and Analysis: Table II indicates that our GNN model associated with the proposed SSHPool can outperform all the alternative methods on four of the five datasets. Only the accuracy of the GNN model associated with the SEPool is a little higher than the proposed SSHPool on the ENZYMES dataset. The effectiveness for our GNN model associated with the proposed SSHPool are threefold. First, all the alternative GNN models or the GNN models associated with different alternative graph pooling operations mainly rely on the node information propagation over the global graph structure. As a result, all these alternative methods suffer the over-smoothing problem, and the node representations extracted by these methods tend to be similar and not discriminative for classification. By contrast, the proposed SSHPool can adaptively decompose the input graph into separated subgraphs, and thus perform the individual local convolution operation on each subgraph structures. Since there is no connection between different separated subgraphs, the convolution operation can only propagate the node information within each subgraph and the node information cannot be propagated between different subgraphs. As a result, the proposed SSHPool can significantly reduce the over-smoothing problem and provide discriminative structural representations for graphs. Second, the WLSK kernel associated with a C-SVM is a typical instance of shallow learning. Moreover, the required kernel computation is also separated from the C-SVM classifier, and cannot provide an end-to-end learning mechanism. By contrast, our GNN model associated with the proposed SSHPool is an end-to-end deep learning approach, naturally having better classification performance. Third, the GNN models associated with the alternative pooling operations may suffer from the degradation problem, since the associated hierarchical pooling mechanisms usually have multiple layer-wise hierarchical structures. By contrast, to overcome the possible degradation problem that maybe caused by the deeper hierarchical structures, we employ an attention layer to accomplishes the interaction between the global features extracted from the SSHPool and the initial embeddings from the original graph, further improving the effectiveness of the proposed GNN-based model.
@@
Finally, we observe that the proposed SSHPool-based GNN model significantly outperforms SSHPool(non) based model, i.e., the proposed SSHPool without the Global Attention Layer. This indicates that our Global Attention Layer can really address the possible degradation problem caused by the deeper architecture of the proposed SSHPool. Moreover, although the SSHPool(non) cannot achieve the best performance due to the possible degradation problem, it still outperforms most of the alternative methods, especially the GNN models associated with other pooling method in terms of the averaged ranking. This again indicates that the separated subgraphs associated with the local graph convolution operations can tremendously improve the effectiveness.
IV-B The Sensitive Analysis with the Depth of the SSHPool
@@
To explore the effectiveness of the proposed SSHPool one step further, we evaluate how the classification performance is influenced by the depth of the SSHPool (i.e., the stacked layers of the SSHPool). Moreover, wee also perform the same experiment with the DiffPool and the SSHPool(non). For the experiments, we utilize the D&D dataset as an instance for these pooling methods, since we will observe the similar phenomenon on other datasets. Specifically, the experimental results are shown in Figure 4. We observe that the proposed SSHPool can outperform the alternative pooling methods associated with all different pooling depth, again demonstrate the effectiveness for the propose SSHPool. Moreover, we observe that the accuracies tend to be lower with the deeper depth, due to the possible degradation problem. However, we find that the accuracies of the proposed SSHPool tend to decrease slowly than the alternative pooling methods, again demonstrating the associated Global Attention Layer for the proposed SSHPool can significantly address the degradation problem.
Figure 4: Sensitivity analysis with the depth of the pooling methods.
@@ V Conclusion
In this paper, we have developed a novel SSHPool method, that cannot only extract effective hierarchical characteristics of graphs but also address the over-smoothing problem through the local separated substructures. Moreover, we have developed a GNN model associated with the proposed SSHPool for graph classification. Experimental results have demonstrated the effectiveness of the proposed SSHPool.

### v2

@@ IV-A Comparisons with Classical GNN and Pooling Methods
Experimental Settings: We evaluate the classification performance of our GNN-based framework associated with the proposed SSHPool operation, and compare it against several alternative methods, including one classical graph kernel associated with C-Support Vector Machine (C-SVM) [34], seven GNN-based models with different pooling methods or optimization approaches. All these baseline models for the evaluation are introduced as follows.
@@
  • SUGAR [19] is an embedding-based GNN model with hierarchical subgraph-level pooling.
  • DiffPool [17] is a GNN model with differentiable graph pooling module that generates representations of coarsened graphs. In the hierarchical architecture, there is an assignment matrix to compress the graph.
@@
For the alternative graph kernels, we search the optimal hyperparameters for the graph kernel on each dataset. For the GNN models or the GNN models associated with the proposed and alternative graph pooling operations, the Adam optimizer is set with the same learning rate of $31e-$ 3 in our experiments, and each method is trained with 100 epochs. Moreover, during the training process of these GNN models, we set the hidden dimension as 128, the dropout as 0.5, and the batch size as 32. In order to compare all these methods fairly, for either the graph kernels or the GNN models we adopt the 10-fold cross validation to compute the classification accuracy, and repeat the experiment for 10 times on each dataset. We show the average classification accuracy ($\pm$ standard error) and the average rank for each method in Table II. Finally, for the proposed SSHPool, we set the largest layer $L$ as $3$, then node numbers of the three layers follow $\{128,32,8\}$ where the assignment ratio $\alpha$ is 0.25.
@@
Datasets D&D PTC PROTEINS NCI1 NCI109 IMDB-B IMDB-M Avg. Rank
WLSK 74.38$\pm$0.69 59.23$\pm$0.45 71.70$\pm$0.67 70.32$\pm$0.40 69.71$\pm$0.52 64.48$\pm$0.90 43.38$\pm$0.75 9.15
DGCNN 75.34$\pm$0.78 58.34$\pm$2.39 73.21$\pm$0.34 67.78$\pm$1.02 67.42$\pm$0.67 67.45$\pm$0.83 46.33$\pm$0.73 9.00
SUGAR 78.62$\pm$0.26 65.72$\pm$2.15 76.81$\pm$0.31 74.93$\pm$1.34 73.72$\pm$0.66 71.43$\pm$0.93 48.12$\pm$1.76 3.86
DIFFPool 77.24$\pm$0.46 63.39$\pm$1.03 74.86$\pm$0.35 62.32$\pm$1.90 61.98$\pm$1.98 70.12$\pm$0.63 47.20$\pm$1.81 8.14
ASAPool 75.75$\pm$0.45 64.71$\pm$2.41 75.53$\pm$0.94 71.48$\pm$0.42 70.07$\pm$0.55 71.25$\pm$0.58 48.67$\pm$1.34 5.86
ABDPool 74.13$\pm$0.52 63.82$\pm$2.37 73.24$\pm$0.91 71.54$\pm$1.28 71.78$\pm$1.35 70.58$\pm$0.71 50.63$\pm$1.47 6.43
SEPool 77.43$\pm$0.23 65.34$\pm$1.03 75.32$\pm$0.34 73.78$\pm$0.31 73.17$\pm$0.42 70.34$\pm$0.53  50.31$\pm$1.72 4.86
AdamGNN 80.12$\pm$0.34 66.73$\pm$0.68 77.04$\pm$0.78 75.67$\pm$0.30 72.99$\pm$0.52 72.85$\pm$0.67 49.35$\pm$0.96 2.86
SSHPool(non) 79.31$\pm$0.36 63.91$\pm$1.35 77.18$\pm$0.43 74.18$\pm$0.45 73.33$\pm$0.38 71.80$\pm$0.39 48.13$\pm$1.03 3.86
SSHPool 81.80$\pm$0.48 67.74$\pm$1.43 79.38$\pm$0.28 75.91$\pm$0.14 74.45$\pm$0.16 73.02$\pm$0.47 51.14$\pm$1.28 1.00
@@
Experimental Results and Analysis: Table II indicates that our GNN model associated with the proposed SSHPool can outperform all the alternative methods on the seven standard datasets. The effectiveness of our GNN model associated with the proposed SSHPool is threefold. First, all the alternative GNN models or the GNN models associated with different alternative graph pooling operations mainly rely on the node information propagation over the global graph structure. As a result, all these alternative methods suffer the over-smoothing problem, and the node representations extracted by these methods tend to be similar and not discriminative for classification. By contrast, our proposed SSHPool can adaptively decompose the input graph into separated subgraphs, and thus perform the individual local convolution operation on each subgraph. Since there is no connection between different separated subgraphs, the convolution operation can only propagate the node information within each subgraph and the node information cannot be propagated between different subgraphs. As a result, the proposed SSHPool can significantly reduce the over-smoothing problem and provide discriminative structural representations for graphs. Second, the WLSK kernel associated with a C-SVM is a typical instance of shallow learning. Moreover, the required kernel computation is also separated from the C-SVM classifier, and cannot provide an end-to-end learning mechanism. By contrast, our GNN model associated with the proposed SSHPool is an end-to-end deep learning approach, naturally having better classification performance. Third, the GNN models associated with the alternative pooling operations may suffer from the degradation problem, since the associated hierarchical pooling mechanisms usually have multiple layer-wise hierarchical structures. By contrast, to overcome the possible degradation problem that may be caused by the deeper hierarchical structures, we employ an attention layer to accomplish the interaction between the global features extracted from the SSHPool and the initial embeddings from the original graph, further improving the effectiveness of the proposed GNN-based model.
@@
Finally, we observe that the proposed SSHPool-based GNN model significantly outperforms the SSHPool(non) based model, i.e., the proposed SSHPool without the Global Attention Layer. This indicates that our Global Attention Layer can really address the possible degradation problem caused by the deeper architecture of the proposed SSHPool. Moreover, although the SSHPool(non) cannot achieve the best performance due to the possible degradation problem, it still outperforms most of the alternative methods, especially the GNN models associated with other pooling methods in terms of the averaged ranking. This again indicates that the separated subgraphs associated with the local graph convolution operations can tremendously improve the effectiveness.
IV-B The Sensitive Analysis
@@
The Depth of the SSHPool. To explore the effectiveness of the proposed SSHPool one step further, we evaluate how the classification performance is influenced by the depth of the SSHPool (i.e., the stacked layers of the SSHPool). Moreover, we also perform the same experiment with the DiffPool and the SSHPool(non). For the experiments, we utilize the D&D dataset as an instance for these pooling methods, since we will observe the similar phenomenon on other datasets. Specifically, the experimental results are shown in Figure 4. We observe that the proposed SSHPool can outperform the alternative pooling methods associated with all different pooling depths, again demonstrating the effectiveness of the proposed SSHPool. Moreover, we observe that the accuracies tend to be lower with the deeper depth, due to the possible degradation problem. However, we find that the accuracies of the proposed SSHPool tend to decrease slowly than the alternative pooling methods, again demonstrating the Global Attention Layer for the proposed SSHPool can significantly address the degradation problem.
Figure 4: Analysis with the depth.
The Assignment Ratio of the SSHPool. To further explore the assignment process of the proposed SSHPool, we evaluate how the classification performance varies associated with different node assignment ratios as $\alpha$. Specifically, we set the ratios as $\alpha\in\{0.5,0.25,0.125\}$. We adopt the D&D, PTC, PROTEINS, and IMDB-B for this evaluation and the experimental results are shown in Figure 5. Overall, the SSHPool associate with the assignment ratio $\alpha=0.25$ outperforms that associated with the assignment ratios $\alpha=0.125$ and $\alpha=0.5$, i.e., the assignment ratios can significantly influence the performance of the proposed SSHPool. This is due to the fact that the assignment ratio determines the number of the separated subgraphs for the next layer of the SSHPool. If the ratio is very small, the nodes of the current layer will be assigned into less clusters, and the size of the resulting coarsened graph from the SSHPool will be quite small than the original graph. On the other hand, if the ratio is very large, the nodes of the current layer will be assigned into more clusters, and the size of the resulting coarsened graph from the SSHPool tends to be as same as the original graph. Both of the above situations cannot represent a fine hierarchical structural information for the SSHPool. This evaluation indicates that the appropriate selection of the assignment ratio is necessary for the proposed SSHPool.
Figure 5: Analysis with the assignment ratio.
@@ V Conclusion
In this paper, we have developed a novel SSHPool method, that cannot only extract effective hierarchical characteristics of graphs but also address the over-smoothing problem through the local separated substructures. Moreover, we have developed a GNN model associated with the proposed SSHPool for graph classification. Experimental results have demonstrated the effectiveness of the proposed SSHPool. Moreover, we also discuss the reasons why our proposed SSHPool is successful in addressing these problems.